Solve this system of equations: x-y=5
______
3x+2y=10

Answer by stanbon(75887)

Solve this system of equations:
x-y=5
3x+2y=10
---------
2x - 2y = 10
3x + 2y = 10
------
Add and solve for "x":
5x = 20
x = 4
-------
Solve for "y":
x - y = 5
4 - y = 5
y = -1
=============
Cheers,
Stan H.
===========